OpenStudy (my4babiesandme):

sqrt 7a + 2sqrt 175a^3

OpenStudy (king):

Is this =0?

OpenStudy (anonymous):

or is it to be factored?

OpenStudy (my4babiesandme):

sqrt 7a + 2sqrt 175a^3= simplify

OpenStudy (anonymous):

if so = sqrt7a + 2* sqrt7 sqrt25 a^3 = sqrt7 a + 10 sqrt7 a^3 = sqrt7a(1 + 10a^2)

OpenStudy (anonymous):

i have assumed that the square roots refer only to the numbers

OpenStudy (my4babiesandme):

i do not know these things confuse me

OpenStudy (anonymous):

probably refer only to numbers otherwise if they referred to letters as well they would have been written as sqrt (7a)

OpenStudy (my4babiesandme):

thanks